Many doors in public places are fitted with simple and inexpensive mechanisms that are designed to close the doors while reduM = 70kg C = 1Ns/m k = 26N/m FO = 5N/S 9 = 95 (duration of the couple in seconds) b = 45N (magnitude of the step in N) z = 2.
Answer #1

Please find attached image for the solution

Gils) = 985+2.4) 15+4:3) open loop transfer function. x(s) G (s) = 45(5+2.4) 15+48)65245+ ) for root locus - G (3) HG) = Kas(Date Page n/2) = 1 n(P) 3 O = (29+1)xido 2 angle of asymptotes = 90 270 centroid = { Real part of O. L. pole-{ real part of o